OpenStudy (anonymous):

HELP PLEASE :) Let R=(rij)24 with rij=i-j, find R

OpenStudy (anonymous):

\[(x_{ij})_{24}, with r _{ij}=i-j\]

OpenStudy (anonymous):

its and r instead of the x

OpenStudy (anonymous):

@jim_thompson5910 @SolomonZelman any idea?

OpenStudy (anonymous):

@Jhannybean

OpenStudy (jhannybean):

\[\large \sf R=(r_{ij})_{24} ~,~ r_{ij}= i-j\] @iambatman @wio help !

OpenStudy (jhannybean):

Hmm..... \[(i-j)_{24}\] I'm not quite sure how to solve these :\

OpenStudy (anonymous):

Can we be more clear here about notation?

OpenStudy (anonymous):

What does the \(24\) mean?

OpenStudy (anonymous):

matrix dimension, 2x4

OpenStudy (anonymous):

\[ R=\begin{bmatrix} r_{1,1}&r_{1,2}&r_{1,3}&r_{1,4}\\ r_{2,1}&r_{2,2}&r_{2,3}&r_{2,4} \end{bmatrix}=\begin{bmatrix} 1-1&1-2&1-3&1-4\\ 2-1&2-2&2-3&2-4 \end{bmatrix} \]

OpenStudy (anonymous):

\[ R==\begin{bmatrix} 0&-1&-2&-3\\ 1&0&-1&-2 \end{bmatrix} \]

OpenStudy (anonymous):

wait how did you get those number ?, not in the solution matrix, the previous one

OpenStudy (anonymous):

oh i got it

OpenStudy (anonymous):

\[ r_{ij} = i-j\implies r_{1,1} = 1-1 \]

OpenStudy (anonymous):

thank you sooo much :)